### v1

@@ 6 The CoomSuite Workbench
The CoomSuite (Baumeister et al., 2025a) is intended to serve as a workbench for experimentation with industrial-scale product configuration problems. While the included benchmark collection can be utilized with other paradigms, its current infrastructure is primarily geared towards ASP. Specifically, the CoomSuite is available as a Python package, installable via pip. It includes a (customizable) ANTLR v4 parser to convert Coom specifications into facts, along with an ASP encoding to harmonize the fact format with the chosen configuration encoding. The current distribution includes four scalable benchmark series, a single ASP encoding covering all essential configuration concepts needed for the three Coom language fragments, as well as an additional one for the hybrid solver fclingo (Cabalar et al., 2025).
@@
For each of the three Coom language fragments, we include a benchmark set in the CoomSuite detailed in Table 1. The two language extensions Coom[p*] and Coom[x*] currently have no dedicated benchmark sets. Additionally, the CoomSuite includes the benchmark set of a Restaurant corresponding to the Coom[x] language. Here, the aim is to configure the assignment of a given number of chairs to tables of different sizes. It makes use of partonomy as well as simple numeric constraints and the scalable factor is the total number of chairs needed.
@@
To provide a glimpse into the usage of the CoomSuite, we conducted sample benchmarks comparing the standard ASP encoding (Section 4.2.2) with that designed for fclingo (Section 4.3), which utilizes integer variables. We ran all instances of the benchmarks in Table 1 on a compute cluster with Intel Xeon E5-2650v4@2.9GHz CPUs with 64GB of memory running Debian Linux 10.444https://www.cs.uni-potsdam.de/bs/research/labs.html#hardware We used a timeout of 300 seconds and limited the memory to 16GB per instance. Figure 8 shows the runtimes of finding one stable model for both solvers clingo and fclingo. On the two non-numeric domains (Core and CityBikeFleet) clingo performs better than fclingo. While clingo is able to solve most instances within a couple seconds (and takes up to 60 seconds for bigger instances), fclingo times out on some of the instances of the Core domain. In general the runtimes of fclingo are about one order of magnitude higher than those of clingo. However, for the TravelBikeFleet, fclingo clearly outperforms clingo which has problems with the (large) numeric ranges. On the contrary, fclingo can handle these effortlessly due to its native handling of integer variables. Note that when clingo times out it is usually during grounding, while for fclingo this usually happens during its internal preprocessing which uses clingo’s Python API.
@@ 7 Discussion
Our work provides not only the first publicly available ASP implementation of Coom but also (indirectly) establishes first semantic underpinnings for Coom. However, this is just the starting point. Future challenges include addressing unimplemented Coom features such as optimization and explanations, as well applying and further developing alternative ASP configuration encodings (Falkner et al., 2015; Gençay et al., 2019; Rühling et al., 2023) within this uniform industrial-scale setting. Concrete improvements include a normalization of constraints during parsing which enables the reutilization of information during solving, studying a more native representation of Coom for fclingo, and the development of a truly incremental approach for unbounded cardinalities using clingo’s multi-shot solving capabilities. While the current approach to the latter provides a simple and general baseline solution, a multi-shot approach will most likely provide a much better performance. However, this requires greater modifications to the ASP encoding and further study on how to handle multiple, especially nested, open cardinalities effectively as well as an extension of the benchmark sets for the sake of an exhaustive evaluation.
As what concerns transparency and user integration, we can leverage ASP-driven visualization (Hahn et al., 2022) and enhance the capabilities for interactive exploration of the configuration space by means of further conceptual and technical development of the UI.

### v2

@@ 6 The CoomSuite Workbench
The CoomSuite (Baumeister et al. 2025a) is intended to serve as a workbench for experimentation with industrial-scale product configuration problems. While the included benchmark collection can be utilized with other paradigms, its current infrastructure is primarily geared towards ASP. Specifically, the CoomSuite is available as a Python package, installable via pip. It includes a (customizable) ANTLR v4 parser to convert Coom specifications into facts, along with an ASP encoding to harmonize the fact format with the chosen configuration encoding. The current distribution includes four scalable benchmark series, a single ASP encoding covering all essential configuration concepts needed for the three Coom language fragments, as well as an additional one for the hybrid solver flingo (Cabalar et al. 2025).
@@
For each of the three Coom language fragments, we include a benchmark set in the CoomSuite detailed in Table 1. The two language extensions Coom[p*] and Coom[x*] currently have no dedicated benchmark sets. Additionally, the CoomSuite includes the benchmark set of a Restaurant corresponding to the Coom[x] language, however, with much simpler numerical calculations. Here, the aim is to configure the assignment of a given number of chairs to tables of different sizes. It makes use of partonomy as well as simple numeric constraints and the scalable factor is the total number of chairs needed. We decided to not show the results of the Restaurant domain here, as they were similar to the Core and CityBikeFleet domains.
@@
To provide a glimpse into the usage of the CoomSuite, we conducted sample benchmarks comparing the standard ASP encoding (Section 4.2.2) with that designed for flingo (Section 4.3), which utilizes integer variables. We ran all instances of the benchmarks in Table 1 on a compute cluster with Intel Xeon E5-2650v4@2.9GHz CPUs with 64GB of memory running Debian Linux 10.666https://www.cs.uni-potsdam.de/bs/research/labs.html#hardware We used a timeout of 300 seconds and limited the memory to 16GB per instance. Figure 8 shows the runtimes of finding one stable model for both solvers clingo and flingo. On the two non-numeric domains (Core and CityBikeFleet) clingo performs better than flingo. While clingo is able to solve most instances within a couple seconds (and takes up to 60 seconds for bigger instances), flingo times out on some of the instances of the Core domain. In general the runtimes of flingo are about one order of magnitude higher than those of clingo. However, for the TravelBikeFleet, flingo clearly outperforms clingo due to the problem of the latter with large numeric ranges. In contrast, flingo can handle these effortlessly due to its native handling of integer variables. Note that when clingo times out it is usually during grounding, while for flingo this usually happens during its internal preprocessing which uses clingo’s Python API.
@@ 7 Discussion
Our work provides not only the first publicly available ASP implementation of Coom but also (indirectly) establishes first semantic underpinnings for Coom. However, this is just the starting point. Future challenges include addressing unimplemented Coom features such as optimization and explanations, as well applying and further developing alternative ASP configuration encodings (Falkner et al. 2015; Gençay et al. 2019; Rühling et al. 2023) within this uniform industrial-scale setting. Concrete improvements include a normalization of constraints during parsing which enables the reutilization of information during solving, studying a more native representation of Coom for flingo, and the development of a truly incremental approach for unbounded cardinalities using clingo’s multi-shot solving capabilities. While the current approach to the latter provides a simple and general baseline solution, a multi-shot approach will most likely provide a much better performance. However, this requires greater modifications to the ASP encoding and further study on how to handle multiple, especially nested, open cardinalities effectively as well as an extension of the benchmark sets for the sake of an exhaustive evaluation.
As what concerns transparency and user integration, we can leverage ASP-driven visualization (Hahn et al. 2022) and enhance the capabilities for interactive exploration of the configuration space by means of further conceptual and technical development of the UI.
@@
Appendix A flingo language
This section is taken and adapted from a forthcoming paper on the system flingo, and will be removed on publication.
We provide here a short introduction to the language of system flingo and illustrate its behavior with examples. The main feature of flingo is that the values assigned to integer variables need to be founded. Recall that in ASP, propositional atoms are false by default and any true propositional atom needs to be founded in the facts of the program by a sequence of rules. Analogously, in flingo, integer variables are undefined by default and any defined integer variable needs to be founded in the facts of the program by using the rules. This behavior is inspired by the Logic of Here-and-There with constraints ($\textrm{HT}_{\!c}$Cabalar et al. (2016)) and is significantly different from other existing CASP solvers such as clingcon. As an example, consider the program
{a}. &sum{x} = 1 :- a. (1)
which has exactly two flingo stable models: one where a is false and x is undefined, and another where a is true and x is assigned the value $1$. Both of these are also clingcon stable models, but clingcon has infinitely many stable models where a is false and x is assigned any integer. The second distinct feature of flingo is the support of aggregates over integer variables, something that is not supported by other existing CASP solvers.
We start by describing flingo’s conditional terms, which are inspired by aggregate elements in ASP and $\textrm{HT}_{\!c}$ conditional terms. A product term is either an integer, an integer variable, or an expression of the form nxn*x where $n$ is an integer and $x$ is an integer variable. A flingo conditional term is an expression of the form
$s:L$ (2)
where $s$ is a product term and $L$ is a comma separated list of propositional literals. The term $s$ is the one considered when the condition $L$ is satisfied, otherwise a default term is used, which is context-dependent and introduced below.
A flingo term is either a product term or a flingo conditional term. The following constraint atoms are supported by flingo:
${\texttt{\&sum}}\{t_{1};\ldots;t_{n}\}\prec s$ (3)
${\texttt{\&sus}}\{t_{1};\ldots;t_{n}\}\prec s$ (4)
${\texttt{\&max}}\{t_{1};\ldots;t_{n}\}\prec s$ (5)
${\texttt{\&min}}\{t_{1};\ldots;t_{n}\}\prec s$ (6)
${\texttt{\&df}}\{x\}$ (7)
where each $t_{i}$ is a flingo term, $s$ is an product term, $x$ is an integer variable, and $\prec$ is one of the following relations <=, =, !=, <, >, >=. The corresponding default terms are defined as 0 for $\mathtt{sum}$ and $\mathtt{sus}$, while $\mathtt{max}$ and $\mathtt{min}$ use the minimum and maximum elements of the integer domain, respectively. The operation $\mathtt{sus}$ stands for strict sum (more precisely $\mathtt{sus}$ stands for sum strict) while $\mathtt{sum}$ is its non-strict version. Intuitively, the non-strict sum discards undefined variables, while the strict does not. If any variable is undefined, the constraint atom with strict sum is false. As an example, consider the constraint atoms
${\texttt{\&sum}}\{x_{1};x_{2}\}\ \mbox{<=}\ 3$ (8)
${\texttt{\&sus}}\{x_{1};x_{2}\}\ \mbox{<=}\ 3$ (9)
and the valuation $v=\{x_{1}\mapsto 1\}$}. In this valuation, variable $x_{2}$ is undefined. Atom (8) is satisfied by valuation $v$ because $v(x_{1})=1\leq 3$ and $x_{2}$ is undefined, and thus is discarded from the sum. In contrast, atom (9) is not satisfied simply because $x_{2}$ is undefined. Operations $\mathtt{max}$ and $\mathtt{min}$ always work in a non-strict way: for instance, $\mathtt{max}$ always return the maximum value among the defined terms in the set.
A flingo atom is either a propositional atom or a constraint atom of the forms (3-7). A flingo literal is any flingo atom $b$, its default negation $\mathtt{not}\ b$ or a flingo atom preceded by two negations $\mathtt{not}\ \mathtt{not}\ b$. A flingo rule is an expression of the form
$a_{0}\ \texttt{:-}\ l_{1},\ldots,l_{n}$ (10)
where $a_{0}$ is a flingo atom or the symbol $\bot$ denoting falsity, and each $l_{i}$ is a flingo literal. Note that, syntactically, every clingcon rule is also a flingo rule.
One may raise the question of why we do not use &sum to denote the strict sum in flingo. The reason is that we aim to keep the syntax of flingo as close as possible to the usual ASP syntax for aggregates, and in ASP, aggregates behave in a non-strict way. For instance, the flingo program
a :- &sum{x : p} = 0. p.
has a single stable model where a and p are true and $x$ is undefined, while if we replace &sum with &sus, the resulting program leaves a undefined. This program can be rewritten in ASP using the aggregate #sum as follows:
a :- #sum{ X: val(X), p } = 0. p.
whose unique stable model is $\{\texttt{a}\}$.
The relation between flingo and clingcon is similar to the relation between the stable models semantics and classical propositional logic, except that here we need to restrict to stable models without undefined integer variables. Since flingo may leave integer variables undefined, a program in flingo may have stable models that the corresponding program in clingcon does not have. As an example, consider again the clingcon program consisting of the single rule $\texttt{a}\ \texttt{:-}\ {\texttt{\&sum}}\{x\}=x$} and its corresponding flingo program consisting of the single rule $\texttt{a}\ \texttt{:-}\ {\texttt{\&sus}}\{x\}=x$}. The first program has infinitely many clingcon answer sets where a is true and $x$ is assigned any integer, while the second program has a single flingo answer set where a is false and $x$ is undefined.
Continuing with the analogy between the stable models semantics and classical propositional logic, in the latter, we can capture all models of a propositional theory by adding choice rules for each propositional variable. Similarly, we can capture all stable models of a clingcon program by adding facts of the form ${\texttt{\&sum}}\{x\}=x$} for each integer variable $x$ in flingo. These facts act as a kind of choice rule for integer variables, allowing them to pick any value.
Assignments
A common pattern in flingo programs is to define a variable in terms of other variables. While constraint atoms (3-7) do not make any difference between the variables occurring on them, an assignment rule allow us to naturally express this pattern. We define an assignment rule as one of the two expressions
${\texttt{\&sus}}\{t_{1};\ldots;t_{n}\}\ \texttt{=:}\ s\ \ \ \texttt{:-}\ \ \ l_{1},\ldots,l_{m}$ (11)
${\texttt{\&in}}\{t_{l}..t_{u}\}\ \texttt{=:}\ s\ \ \ \texttt{:-}\ \ \ l_{1},\ldots,l_{m}$ (12)
where each $t_{i}$ and $t_{l}$, $t_{u}$ are flingo terms, $s$ is an product term, and each $l_{i}$ is a flingo literal. The operator $\mathtt{in}$ functions like a choice rule with a fixed range for integer variables, assigning a value in the range of $t_{l}$ and $t_{u}$ to $s$.
An assignment rule (11) or (12) only defines the variables in $s$ so that the constraint atom obtained by replacing the assignment operator =: with equality = is satisfied. However, it does not define any variables in $t_{1},\ldots,t_{n}$. For instance, the assignment rule &sus{x ; y} =: z ensures that z is defined as the sum of x and y, when both x and y are defined. However, if x or y are undefined, then the rule has no effect on the value assigned to z. As an example, if we add facts x = 1 and y = 2 to the above assignment, then the resulting program has a single stable model where x, y, and z are assigned $1$, 22, and $3$, respectively. If instead we add only the fact x = 1, then the resulting program has a single stable model where x is assigned $1$, while y and z are undefined. And if instead we add the facts x = 1 and z = 5, then the resulting program has a single stable model where x and z are assigned $1$ and $5$, respectively, while y is undefined.
An assignment rule of the form of (11) or (12) is an abbreviation, respectively, for the rules:
${\texttt{\&sus}}\{t_{1};\ldots;t_{n}\}\ \mbox{=}\ s\ \ \texttt{:-}\ \ \ l_{1},\ldots,l_{m},{\texttt{\&df}}(x_{1}),\dotsc,{\texttt{\&df}}(x_{k})$ (13)
${\texttt{\&in}}\{t_{l}..t_{u}\}\ \mbox{=}\ s\ \ \ \texttt{:-}\ \ \ l_{1},\ldots,l_{m},{\texttt{\&df}}(x_{1}),\dotsc,{\texttt{\&df}}(x_{k})$ (14)
where $x_{1},\ldots,x_{k}$ are all the the integer variables occurring in $t_{1},\dotsc,t_{n}$ or $t_{l}$ and $t_{u}$, respectively.